# Spatial Association Between Near-Misses and Accident Blackspots in Sydney, Australia: A Getis-Ord $G_{i}^{*}$ Analysis

### v1

@@ Abstract
Road safety management teams utilize on historical accident logs to identify blackspots, which are inherently rare and sparse in space and time. Near-miss events captured through vehicle telematics and transmitted in real-time by connected vehicles reveal a unique potential of prevention due to their high frequency nature and driving engagement on the road. There is currently a lack of understanding of the high potential of near-miss data in real-time to proactively detect potential risky driving areas, in advance of a fatal collision. This paper aims to spatially identify clusters of reported accidents (A) versus high-severity near-misses (High-G) within an urban environment (Sydney, Australia) and showcase how the presence of near-misses can significantly lead to future crashes in identified risky hotspots. First, by utilizing a 400m grid framework, we identify significant crash hotspots using the Getis-Ord $G_{i}^{*}$ statistical approach. Second, we employ a Bivariate Local Moran’s I (LISA) approach to assess and map the spatial concordance and discordance between official crash counts (A) and High-G counts from nearmiss data (High-G). Third, we classify areas based on their joint spatial patterns into: a) High-High (HH) as the most riskiest areas in both historical logs and nearmiss events, High-Low (HL) for high crash logs but low nearmiss records, c) Low-High (LH) for low past crash records but high nearmiss events, and d) Low-Low (LL) for safe areas. Finally, we run a feature importance ranking on all area patterns by using a contextual Point of Interest (POI) count features and we showcase which factors are the most critical to the occurrence of crash blackspots.
@@
connected vehicles, nearmisses, blackspots, statistical inference
@@ II Related Works
These limitations spurred a significant shift towards proactive methodologies leveraging Surrogate Safety Measures (SSMs) [12]. The foundation lies in the Traffic Conflict Technique (TCT), formalized decades ago to systematically observe near-miss events [3, 14]. A near-miss or traffic conflict is generally defined as an interaction necessitating an evasive maneuver to avoid a collision [12, 3]. These non-crash events occur far more frequently than actual crashes, providing statistically richer datasets for analyzing underlying traffic risks and evaluating countermeasures without relying on sparse crash data [12, 4]. Consequently, research has focused on developing and applying various SSM indicators derived from detailed observational or sensor data. Common indicators include temporal measures like Time-to-Collision (TTC) and Post-Encroachment Time (PET), deceleration requirements like DRAC, and kinematic indicators such as speed, lateral deviation, or harsh events (e.g., rapid braking/acceleration) [12, 4]. Data for SSM calculation is increasingly sourced from video analytics platforms employing computer vision and AI [15, 16, 17], instrumented vehicles in Naturalistic Driving Studies (NDS) [2, 18], smartphone sensors [4], and Connected Vehicle (CV) data streams (e.g., Basic Safety Messages) [4, 19].
Analyzing crash causation to develop effective safety countermeasures is essential, yet hindered by the low frequency of actual crash events, particularly within rich datasets from Naturalistic Driving Studies (NDS) [20]. NDS provides detailed real-world driving data but typically captures few crashes relative to the volume of driving. This data scarcity necessitates the use of Surrogate Safety Measures (SSMs) – observable events, like traffic conflicts or near-crashes, thought to be correlated with crash risk [4]. However, establishing the validity of SSMs as reliable proxies for crash risk remains a significant and ongoing challenge in traffic safety research [21].
Near-crashes, commonly defined as events requiring a rapid evasive maneuver to avoid a collision, are frequently used SSMs, especially in NDS analysis. Their application often relies on the “causal continuum” hypothesis, which posits that near-crashes and crashes arise from largely similar or identical causal factors. This assumption is considered plausible and supported by reported correlations between conflict/near-crash frequency and historical crash rates [4]. Foundational research by Guo et al. [22] provided critical support for this approach by demonstrating that near-crashes often exhibit kinematic signatures similar to crashes and appear to share common underlying causal mechanisms or contributing factors. These findings established near-crashes as viable and effective surrogates, allowing researchers to analyze the more abundant near-crash data to understand risk and causation.
@@
While near-miss frequency indicates the prevalence of conflicts, metrics reflecting near-miss severity, such as high G-force events (often denoted NM+G) captured by inertial sensors in vehicles or smartphones, offer outlook into the potential consequence or danger level of these interactions [4]. High G-force readings signify rapid changes in velocity indicative of harsh braking or abrupt evasive manoeuvres, which may be associated with more dangerous situations compared to less severe conflicts. However, research directly investigating the relationship between the severity of surrogate events and the severity of actual crashes is less developed than frequency-based comparisons. Specifically, there is limited work that directly compares the spatial patterns of hotspots derived from near-miss severity indicators (like NM+G) with hotspots identified using crash severity data. This study aims to address this specific gap by employing established spatial statistical methods (Getis-Ord $G_{i}^{*}$) to explicitly examine the concordance and discordance between hotspots identified using high G-force near-miss data and those identified using historical crash data, stratified by crash severity. This comparison seeks to clarify the utility of severity-based near-miss data as a spatial proxy for realized accident risk, particularly concerning more severe crash outcomes.
@@ III Case study
  • Points of Interest (POIs): Geographic locations of various amenities, shops, and other features extracted from OpenStreetMap (OSM) data to provide an environmental context.
IV Statistical Analysis Methodology
The core of this study involved a quantitative spatial analysis of reported crashes (A, $n=$3638$$) and high G-force events (High-G, $n=$23999$$) within Sydney (mapped for year 2022), using data aggregated onto a uniform 400m grid. The methodology focused on identifying statistically significant crash hotspots, analyzing the local spatial correlation between crashes and High-G events, and characterizing the resulting spatial patterns using Point of Interest (POI) data. Key stages that have been applied are the following:
IV-A Spatial Framework and Data Aggregation
To facilitate our area-based analysis, a uniform 400m x 400m grid was established across the study region (Projected CRS: EPSG:32756), resulting in $38824$ cells. The size of the grid has been selected after conducting a sensitivity analysis of the best grid size to provide granular view into risky manoeuvres occurring even across smaller secondary roads/nieghbourhood alleys. Data of the reported crashes and High-G events were spatially aggregated onto this grid, yielding cell-level counts for each variable (crash_count, highg_count). These aggregated counts formed the primary input for the spatial statistical analyses described below. Point-based POI data was processed separately for subsequent feature generation (see IV-C).
IV-B Crash Hotspot Identification and Bivariate Correlation
@@
  • Bivariate Spatial Correlation (LISA): Bivariate Local Moran’s I (LISA) analysis was employed to quantify and map the local spatial correlation between crash_count and highg_count. This identified cells exhibiting statistically significant ($p<0.05$) spatial patterns: High Crash-High HighG (HH), Low Crash-Low HighG (LL), High Crash-Low HighG (HL), and Low Crash-High HighG (LH), revealing areas of concordance and discordance between the two indicators.
IV-C Characterization of LISA Patterns using POI Features
Following the identification of distinct spatial correlation patterns via LISA, a characterization focused on the environmental context provided by Points of Interest: POI data (sydney_pois_filtered.csv), containing point locations and types, was spatially joined with the analysis grid. For each grid cell, the number of POIs of each distinct type falling within its boundary was counted using geopandas, generating POI count features (e.g., poi_type_Park, poi_type_School).
IV-D Spatial Clustering of Crash Events
The Getis-Ord $G_{i}^{*}$ statistic was employed to identify statistically significant spatial clusters of reported crashes (A) across the 400m grid cells covering the Sydney study area (circa 2022). This identifies localized concentrations significantly higher (hotspots) or lower (coldspots) than expected given the overall spatial distribution. The resulting spatial pattern of crash hotspots and coldspots is illustrated in Fig. 1.
@@ IV-D Spatial Clustering of Crash Events
Local Spatial Correlation between Crash Counts and High-G Event Counts: To investigate the local relationship between the frequency of reported crashes and the frequency of high G-force (High-G) events, a Bivariate Local Moran’s I (LISA) analysis was performed on the 400m grid cell data. This analysis assesses whether the spatial pattern of crash counts in a given cell is significantly correlated with the spatial pattern of High-G event counts in its neighborhood (using Queen contiguity). The geographical distribution of these local spatial correlations is depicted in 2.
Figure 2: Map of Bivariate Local Moran’s I (LISA) results on the 400m grid, showing the spatial correlation between crash counts and High-G event counts per cell. Colors represent the type of statistically significant ($p<0.05$) local correlation: High-High (Red: high crashes, high High-G), Low-Low (Blue: low crashes, low High-G), High-Low (Pink: high crashes, low High-G), Low-High (Light Blue: low crashes, high High-G). Grey areas indicate no significant local spatial correlation. (Note: Analysis log saved this plot as crash_highg_correlation.pdf)
@@ V Results
V-A Spatial Analysis of Crash and High-G Event Clustering (400m Grid)
The spatial distribution and correlation of reported crashes ($n=$3638$$) and high G-force (High-G) events ($n=$23999$$) were analyzed using a 400m x 400m square grid resolution, encompassing $38824$ cells across the Sydney study area. Initial analysis using Global Moran’s I confirmed significant positive spatial autocorrelation (clustering) for both crash counts (Moran’s I = $0.2080$, $p=0.001$) and High-G event counts (Moran’s I = $0.2586$, $p=0.001$). This indicates an overall tendency for both phenomena to cluster spatially rather than being randomly distributed.
@@ V-A Spatial Analysis of Crash and High-G Event Clustering (400m Grid)
To explore the specific local patterns of spatial association between crash frequency and the near-miss proxy (High-G events), Bivariate Local Moran’s I (LISA) was used (using $p<0.05$ for significance). This method classified the grid cells based on their significant local spatial relationship patterns.
V-B Identification and Classification of Near-Miss Hotspots
This analysis investigates the spatial relationship between potential traffic conflicts, represented by high G-force events (High-G, $n=23,999$), and realized harm, represented by reported crash locations (Crashes, $n=3,638$). The study area was divided into a regular grid with 400m x 400m cells.
To understand the spatial concordance or discordance between these indicators, a Bivariate Local Moran’s I (LISA) analysis was conducted. This spatially correlates the counts of High-G events with the counts of reported Crashes within the same grid cell system (using Queen contiguity weights). The analysis identifies locations where the local concentration of High-G events significantly aligns (or fails to align) with the local concentration of Crashes. Based on the LISA results (significant at $p<0.05$), grid cells were classified into distinct spatial relationship categories. These classifications reveal distinct spatial patterns based on the Bivariate Local Moran’s I (LISA) analysis comparing crash counts and high G-force (High-G) event counts within 400m grid cells (Table I):
@@ V-B Identification and Classification of Near-Miss Hotspots
  • HL (High Crash-Low High-G): Represents 91 grid cells with historically high crash counts that are *not* statistically associated with high counts of the measured High-G events in their vicinity. This suggests that crash risk in these locations might be driven by factors other than those frequently captured by the High-G metric (e.g., specific static hazards, complex intersection designs not inducing frequent harsh braking/swerving, reporting biases, or different types of risky behavior).
  • LH (Low Crash-High High-G): Represents a substantial number of areas (2681 grid cells) characterized by statistically significant high counts of potential conflicts (High-G events) but low counts of actual reported crashes. These ’Emerging Risk’ or ’Near-Miss Hotspot’ locations are of particular interest. They suggest the presence of frequent risky maneuvers or situations, but perhaps mitigating factors (e.g., effective road design allowing recovery, lower speeds, successful evasive actions, under-reporting of minor crashes) currently prevent these from translating into high numbers of reported crashes. These areas warrant proactive monitoring.
@@
These local patterns highlight specific areas of concordance (HH) and discordance (HL, LH) between historical crash data and the High-G near-miss indicator.
TABLE I: Classification and Size of Grid Cells Based on Bivariate Spatial Correlation (LISA, $p<0.05$) between Crash Counts and High G-Force Event Counts.
@@
Comparing these different spatial clusters, particularly the high number of LH cells (2681) versus HH cells (825), suggests that while High-G events are spatially associated with crashes in many hotspots (HH), there are even more areas where frequent near-misses occur without a corresponding high crash history (LH). This highlights the potential of High-G data for proactive safety analysis, identifying areas of concern before they become crash blackspots, and potentially revealing locations where safety interventions or specific road characteristics are effectively mitigating crash outcomes despite frequent risky events. Further analysis of the characteristics differentiating HH, HL, and LH areas is crucial.
Interpretation of Mann-Whitney U Test Results
@@
It is important to note that the Points of Interest (POIs) analyzed in this study originate from two primary sources: OpenStreetMap (OSM) data (specifically using tags like amenity, shop, tourism, etc) and a dedicated traffic light dataset provided by Transport for NSW.
@@
    POIs More Prevalent in High Crash-High HighG Areas: A notable number of POI types were found to be significantly more common in the High-High cluster areas. These include:
    • Retail and Commercial Services: A variety of retail outlets (appliance, bed, radiotechnics, flooring, trophy, lighting, baby_goods, fabric, health_food, pawnbroker, carpet) showed significantly higher mean counts in High-High areas. For many of these, the mean count in Low-Low areas was zero or near-zero, suggesting their presence is a stronger characteristic of the High-High areas within this dataset.
    • Public Amenities and Features: Features such as viewpoint, waste_basket, and bench were also significantly more prevalent in High-High areas. The higher density of benches and viewpoints might suggest areas with higher pedestrian activity or specific urban design features.
    • Educational Facilities: prep_school counts were significantly higher in the High-High areas.
@@
    POIs Less Prevalent in High Crash-High HighG Areas: Perhaps the most striking finding in this category is the significantly lower prevalence of traffic_signals in High-High areas compared to Low-Low areas (p=0.017). This suggests that the High-High areas, despite having more crashes and High-G events, might be less characterized by major, signalized intersections compared to the Low-Low areas used in this comparison. This could point towards differences in road network hierarchy, traffic control strategies, or potentially higher prevalence of unsignalized intersections or different road types (e.g., mid-block segments) contributing to the High-High classification.
@@
    No Significant Difference: Several POI types, including dive_centre, music_school, safety_equipment, grave_yard, arts_centre, handwashing, and tea, did not show a statistically significant difference between the two groups. These POIs were generally rare in both area types, as indicated by their low mean counts.

### v2

@@ Abstract
Conventional road safety management is inherently reactive, relying on analysis of sparse and lagged historical crash data to identify hazardous locations, or crash blackspots. The proliferation of vehicle telematics presents an opportunity for a paradigm shift towards proactive safety, using high-frequency, high-resolution near-miss data as a leading indicator of crash risk. This paper presents a spatial-statistical framework to systematically analyze the concordance and discordance between official crash records and near-miss events within urban environment. A Getis-Ord statistic is first applied to both reported crashes and near-miss events to identify statistically significant local clusters of each type. Subsequently, Bivariate Local Moran’s I assesses spatial relationships between crash counts and High-G event counts, classifying grid cells into distinct profiles: High-High (coincident risk), High-Low and Low-High. Our analysis reveals significant amount of Low-Crash, High-Near-Miss clusters representing high-risk areas that remain unobservable when relying solely on historical crash data. Feature importance analysis is performed using contextual Point of Interest data to identify the different infrastructure factors that characterize difference between spatial clusters. The results provide a data-driven methodology for transport authorities to transition from a reactive to a proactive safety management strategy, allowing targeted interventions before severe crashes occur.
@@
connected vehicles, near-misses, blackspots, statistical inference
@@ II Related Works
These limitations inspire a significant shift towards proactive methodologies using Surrogate Safety Measures (SSMs) [12]. The foundation lies in the Traffic Conflict Technique (TCT) to systematically observe near-miss events [3, 14]. A near-miss or traffic conflict is generally defined as an interaction necessitating an evasive maneuver to avoid a collision [12, 3]. These non-crash events occur far more frequently than actual crashes, providing statistically richer datasets for analyzing underlying traffic risks and evaluating countermeasures without relying on sparse crash data [12, 4]. Consequently, research has focused on developing and applying various SSM indicators derived from detailed observational or sensor data. Common indicators include temporal measures like Time-to-Collision (TTC) and Post-Encroachment Time (PET), deceleration requirements like DRAC, and kinematic indicators such as speed, lateral deviation, or harsh events (e.g., rapid braking/acceleration) [12, 4]. Data for SSM calculation is usually sourced from video analytics platforms using computer vision and AI [15, 16, 17], instrumented vehicles in Naturalistic Driving Studies (NDS) [2, 18], smartphone sensors [4], and Connected Vehicle (CV) data streams [4, 19].
Analyzing crash causation to develop effective safety countermeasures is essential, yet hindered by the low frequency of actual crash events, particularly within rich datasets from Naturalistic Driving Studies (NDS) [20]. NDS provides detailed real-world driving data but typically captures fewer crashes relative to the volume of driving. This data scarcity necessitates the use of SSMs - observable events, like traffic conflicts or near-crashes, thought to be correlated with crash risk [4]. However, establishing the validity of SSMs as reliable proxies for crash risk remains a significant and ongoing challenge in traffic safety research [21].
Near-crashes, commonly defined as events requiring a rapid evasive maneuver to avoid a collision, are frequently used SSMs, especially in NDS analysis. Their application often relies on the “causal continuum” hypothesis, which states that near-crashes and crashes arise from largely similar or identical causal factors. This assumption is considered plausible and supported by reported correlations between conflict/near-crash frequency and historical crash rates [4]. Foundational research by Guo et al. [22] provided critical support for this approach by demonstrating that near-crashes often exhibit kinematic signatures similar to crashes and appear to share common underlying causal mechanisms or contributing factors. These findings established near-crashes as viable and effective surrogates, allowing researchers to analyze the more abundant near-crash data to understand risk and causation.
@@
While near-miss frequency indicates the prevalence of conflicts, metrics reflecting near-miss severity, such as high G-force events (denoted as NM+G) captured by inertial sensors in vehicles or smartphones, offer outlook into the potential consequence or danger level of these interactions [4]. The GforceG-force values for emergency braking vary significantly across multiple studies [25] with deceleration rates of vehicles found to be ranged from 0.49 m/s2 to 8.76 m/s2 with a total weighted average of 2.82 m/s2. In practice, some Automated Emergency Braking systems consider the threshold of 0.47g requiring automated intervention to prevent a collision: Bendix system uses 0.47g as the critical threshold to determine if a driver’s braking is insufficient [26]. The fleet telematics solution Geotab uses 0.47g as a pre-defined setting for flagging harsh cornering events [27]. The threshold of 0.47g is selected to select near miss trajectories for this study.
High G-force readings, which signify rapid deceleration or acceleration, capture the severe vehicle maneuvers (near-miss events). Consequently, they have the potential to act as a robust proxy for identifying locations with a high risk of future collisions. However, research directly investigating the relationship between the severity of surrogate events and the severity of actual crashes is less developed than frequency-based comparisons. Specifically, there is limited work that directly compares the spatial patterns of hotspots derived from near-miss severity indicators (like NM+G) with hotspots identified using crash data. This study aims to address this specific gap by using established spatial statistical methods (Getis-Ord $G_{i}^{*}$) to explicitly examine the concordance and discordance between hotspots identified using high G-force near-miss data and those identified using historical crash data, stratified by crash severity. This comparison seeks to clarify the utility of near-miss data as a spatial proxy for accident risk.
@@ III Case study
  • Points of Interest (POIs): Geographic locations of various amenities, shops, and other features extracted from OpenStreetMap (OSM) data.
IV Methodology
The instantaneous G-force quantifies the acceleration experienced relative to standard gravitational acceleration $g=9.806\,\mathrm{m/s}^{2}$. Let $\mathbf{a}(t)=[a_{x}(t),a_{y}(t),a_{z}(t)]^{\top}\in\mathbb{R}^{3}$ denote the time-dependent vector of longitudinal ($a_{x}$), lateral ($a_{y}$), and vertical ($a_{z}$) accelerations measured by inertial sensors. The Euclidean norm $\left\|\cdot\right\|$ is computed as:
$\left\|\mathbf{a}(t)\right\|=\sqrt{a_{x}(t)^{2}+a_{y}(t)^{2}+a_{z}(t)^{2}}.$
The G-force at time $t$ is defined via the ratio:
$\text{G-force}(t):=\frac{\left\|\mathbf{a}(t)\right\|}{g}=\frac{\sqrt{a_{x}(t)^{2}+a_{y}(t)^{2}+a_{z}(t)^{2}}}{9.806}.$
The trajectory of both the the steering (lateral deviations) and braking (longitudinal deceleration) maneuvers exhibit localized extrema of G force (Fig. 1). Since both types of events have peak kinematic extrema localized within narrow time window, near miss severity can be spatially represented through discrete point of maximum G force within near miss trajectory.
While metrics like Time-to-Collision (TTC) or Post-Encroachment Time (PET) focus on temporal proximity during observed conflicts, this study prioritizes G-force as a kinematic proxy due to critical data limitations. With connected vehicle coverage limited to a small fraction of Sydney’s traffic flow and no access to multi-vehicle collision trajectories, our data is restricted to near-miss events from single vehicles. Furthermore, TTC and PET, as metrics of interaction, require the determination of a conflict point, which is not observable from single-vehicle data.
The core of this study involved a quantitative spatial analysis of reported crashes (A, n=3638n=\$3638\$) and high G-force events (High-G, n=24137n=\$24137\$) within Sydney (mapped for year 2022), using data aggregated onto a uniform 400m grid. The methodology focused on identifying statistically significant crash hotspots, analyzing the local spatial correlation between crashes and High-G events, and characterizing the resulting spatial patterns using POI data. Key stages that have been applied are the following:
Figure 1: Vehicle trajectory segments represented with G-force profiles (x-axis: relative time centered around point of max G force; y-axis: acceleration in $g$)
To facilitate our area-based analysis, a uniform 400m x 400m grid was established across the study region (Projected CRS: EPSG:32756), resulting in $38824$ cells. The size of the grid has been selected after conducting a sensitivity analysis of the best grid size to provide granular view into risky maneuvers. Data of the reported crashes and High-G events were spatially aggregated onto this grid, yielding cell-level counts for each variable (crash_count, highg_count). These aggregated counts formed the primary input for the spatial statistical analyses described below.
@@ IV Methodology
  • Bivariate Spatial Correlation (LISA): LISA analysis was used to quantify and map the local spatial correlation between crash_count and highg_count. This identified cells exhibiting statistically significant ($p<0.05$) spatial patterns: High Crash-High HighG (HH), Low Crash-Low HighG (LL), High Crash-Low HighG (HL), and Low Crash-High HighG (LH), revealing areas of concordance and discordance between the two indicators.
Following the identification of distinct spatial correlation patterns via LISA, a characterization focused on the environmental context provided by Points of Interest.
POIs joined with the spatial grid in this study originate from two primary sources: OpenStreetMap (OSM) amenity data and a dedicated traffic light dataset provided by Transport for NSW.
The Getis-Ord $G_{i}^{*}$ statistic was used to identify statistically significant spatial clusters of reported crashes (A) across the 400m grid cells covering the Sydney study area (circa 2022). This identifies localized concentrations significantly higher (hotspots) or lower (coldspots) than expected given the overall spatial distribution. The resulting spatial pattern of crash hotspots and coldspots is illustrated in Fig. 2.
Getis-Ord $G_{i}^{*}$ excels at identifying statistically significant local clusters of hot or cold spots precisely because it explicitly accounts for spatial autocorrelation by considering the values of neighboring features. In contrast, Kernel Density Estimation (KDE) primarily focuses on estimating the spatial density of point locations and typically assumes an isotropic (uniform in all directions) spread, without accounting for spatial autocorrelation.
@@
Local Spatial Correlation between Crash Counts and High-G Event Counts: To investigate the local relationship between the frequency of reported crashes and the frequency of high G-force (High-G) events, a LISA analysis was performed on the 400m grid cell data. This analysis assesses whether the spatial pattern of crash counts in a given cell is significantly correlated with the spatial pattern of High-G event counts in its neighborhood (using Queen contiguity). The geographical distribution of these local spatial correlations is depicted in 3.
Figure 3: Map of LISA results on the 400m grid, showing the spatial correlation between crash counts and High-G event counts per cell. Colors represent the type of statistically significant ($p<0.05$) local correlation: High-High (Red: high crashes, high High-G), Low-Low (Blue: low crashes, low High-G), High-Low (Pink: high crashes, low High-G), Low-High (Light Blue: low crashes, high High-G). Grey areas indicate no significant local spatial correlation. (Note: Analysis log saved this plot as crash_highg_correlation.pdf)
@@ V Results
The spatial distribution and correlation of reported crashes (n=3638n=\$3638\$) and high G-force (High-G) events (n=24137n=\$24137\$) were analyzed using a 400m x 400m square grid resolution, encompassing $38824$ cells across the Sydney study area. Initial analysis using Global Moran’s I confirmed significant positive spatial autocorrelation (clustering) for both crash counts (Moran’s I = $0.2080$, $p=0.001$) and High-G event counts (Moran’s I = $0.2586$, $p=0.001$). This indicates an overall tendency for both phenomena to cluster spatially rather than being randomly distributed.
@@
V-A Identification and Classification of Near-Miss Hotspots
This analysis investigates the spatial relationship between potential traffic conflicts, represented by high G-force events (High-G, $n=24,137$), and realized harm, represented by reported crash locations (Crashes, $n=3,638$). The study area was divided into a regular grid with 400m x 400m cells.
To understand the spatial concordance or discordance between these indicators, a LISA analysis was conducted. This spatially correlates the counts of High-G events with the counts of reported Crashes within the same grid cell system (using Queen contiguity weights). The analysis identifies locations where the local concentration of High-G events significantly aligns (or fails to align) with the local concentration of Crashes. Based on the LISA results (significant at $p<0.05$), grid cells were classified into distinct spatial relationship categories. These classifications reveal distinct spatial patterns comparing crash counts and high G-force (High-G) event counts within 400m grid cells (Table I):
@@ V-A Identification and Classification of Near-Miss Hotspots
  • HL (High Crash-Low High-G): Represents 91 grid cells with historically high crash counts that are not statistically associated with high counts of the measured High-G events in their vicinity. This suggests that crash risk in these locations might be driven by factors other than those frequently captured by the High-G metric (e.g., specific static hazards, safety-focused intersection designs).
  • LH (Low Crash-High High-G): Represents a substantial number of areas (2681 grid cells) characterized by statistically significant high counts of potential conflicts (High-G events) but low counts of actual reported crashes. These locations with frequent near-misses but low crash history warrant proactive monitoring. They suggest the presence of frequent risky maneuvers or situations, but perhaps mitigating factors (e.g., effective road design allowing recovery, lower speeds, successful evasive actions, under-reporting of minor crashes) currently prevent these from translating into high numbers of reported crashes. These areas warrant proactive monitoring.
@@
TABLE I: Classification and Size of Grid Cells Based on Bivariate Spatial Correlation (LISA, p ¡ 0.05) between Crash Counts and High G-Force Event Counts.
@@
These local patterns highlight specific areas of concordance (HH) and discordance (HL, LH) between historical crash data and the High-G near-miss indicator. Comparing these different spatial clusters, particularly the high number of LH cells (2681) versus HH cells (825), suggests that while High-G events are spatially associated with crashes in many hotspots (HH), there are even more areas where frequent near-misses occur without a corresponding high crash history (LH). This highlights the potential of High-G data for proactive safety analysis, identifying areas of concern before they become crash blackspots, and potentially revealing locations where safety interventions or specific road characteristics are effectively mitigating crash outcomes despite frequent risky events. Further analysis of the characteristics differentiating HH, HL, and LH areas is crucial.
@@
    POIs More Prevalent in High Crash-High HighG Areas: A notable number of POI types were found to be significantly more common in the High-High cluster areas: retail and commercial services (e.g., appliances, lighting), public amenities (benches, viewpoints), and educational facilities. The higher density of benches and viewpoints might suggest areas with higher pedestrian activity.
@@
    POIs Less Prevalent in High Crash-High HighG Areas: Perhaps the most striking finding in this category is the significantly lower prevalence of traffic signals in High-High areas compared to Low-Low areas (p=0.017). This suggests that the High-High areas, despite having more crashes and High-G events, might be less characterized by major, signalized intersections compared to the Low-Low areas used in this comparison. This could point towards differences in road network hierarchy or potentially higher prevalence of unsignalized intersections or different road types (e.g., mid-block segments) contributing to the High-High classification.
@@
    Several types of points of interest, including music schools, safety equipment suppliers, arts centers and tea shops, did not show a statistically significant difference between the two groups. These points of interest were generally rare in both area types, as indicated by their low average counts.
Infrastructure improvements (e.g., signal timing optimization, pedestrian crosswalks) should be prioritized in HH areas given their dual risk of crashes and near-misses. For LH areas, extending monitoring systems for proactive conflict detection, as these sites may evolve into crash blackspots.